A man makes three investments totaling $24,000 with interest rates at 6%, 7%, and 8%. The total annual income is $1720, and the income from 7% investment is $40 less than the combined income from the other investments. Find the amount of each investment.
Answer by richwmiller(17219)

a+b+c=24000,
.06a+.07*b+.08c=1720,
.06a-.07*b+.08c=40
.06a-.07*b-.08c=40
1,1,1,24000
.06,.07,.08,1720
.06,-.07,.08,40
4000,12000,8000